I've been working on the equation: 2x-3y=24 2x+y=8. I find x when i do the equation my way, and the way i've been shown on this site, but i end up with -5y=16, and i get stucc there.
Answer by rcmcc(152)

2x-3y=24
2x+y=8
to solve this question you must opperate the substitution method the second eqation is easier to simplify and isolate y.
2x+y=8
subtract 2x to isolate y
y=8-2x
now substitute in (8-2x) for y in the first equation
2x-3(8-2x)=24
you have to expand the -3 into the brackets so you get the equation
2x-24+6x=24
add 24 to each side and simplify the equation
8x=48
divide both sides by 8 to isolate x
x=6
to find y substitute 6 in for x to any of the above 2 equations, i recomend the second one as it is easier
2x+y=8
2(6)+y=8
12+y=8
subtract 12 from each side
y=-4
so in this equation you should have calculated that x=6 and y=(-4)